OpenStudy (anonymous):

Convert 120 degrees to radian measure in terms of  .

OpenStudy (jmcs1985):

Problems like these make me think of the relationship \[\pi = 180\] So....\[\frac{ \pi }{ 180 } \times \frac{120}{1} = \frac{120\pi}{180}\] Now, divide both numerator and denominator by a common factor - I'm choosing 60: \[\frac{ 2\pi }{ 3 }\] will be your final answer. Hope this helps!

OpenStudy (anonymous):

right

OpenStudy (jmcs1985):

@dannymenz - Did it help clear up the subject matter?

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

If f(x) =3x+4,find f^-1(x)

OpenStudy (aaronq):

ask math questions in the math section - this is biology.